Given x=-9y^2+-5y+-5 find the directrix
Answer #1

Silin) bivan 5 x=-942-sy-s ay?+s4= 0-5 ghts y -x-5 9 ca o 9 ytfalen - 25=- are the day 18 ola 324 2 ts 18 Iss 324 9 lyt j =-*= dn 우 p 3. The directrix in у -\SS 4 5 3 34 hsi- =10 ЗА - - 4 13 » 15 ч• E11 2